# LLM-assisted Explicit and Implicit Multi-interest Learning Framework for Sequential Recommendation

### v1

@@ Abstract.
Multi-interest modeling in current recommender systems (RS) is mainly based on user behavioral data, capturing user interest preferences from multiple dimensions. However, since behavioral data is implicit and often highly sparse, it is challenging to understand users’ complex and diverse interests. Recent studies have shown that the rich semantic information in the text can effectively supplement the deficiencies of behavioral data and provide a new perspective for building more accurate user portraits. Despite this, it is still difficult for small models to directly extract semantic features associated with users’ deep interests. That is, how to effectively align semantics with behavioral information to form a more comprehensive and accurate understanding of user interests has become a critical research problem.
To address this, we propose an LLM-assisted explicit and implicit multi-interest learning framework (named EIMF) to model user interests on two levels: behavior and semantics. The framework consists of two parts: Implicit Behavioral Interest Module (IBIM) and Explicit Semantic Interest Module (ESIM). The traditional multi-interest RS model in IBIM can learn users’ implicit behavioral interests from interactions with items. In ESIM, we first adopt a clustering algorithm to select typical samples and design a prompting strategy on LLM to obtain explicit semantic interests. Furthermore, in the training phase, the semantic interests of typical samples can enhance the representation learning of behavioral interests based on the multi-task learning on semantic prediction and modality alignment. Therefore, in the inference stage, accurate recommendations can be achieved with only the user’s behavioral data. Extensive experiments on real-world datasets demonstrate the effectiveness of the proposed EIMF framework, which effectively and efficiently combines small models with LLM to improve the accuracy of multi-interest modeling.
Recommender System; Sequential Recommendation; Semantic Alignment; Multi-Interest Learning; Large Language Models
submissionid: 353copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RS) (Bobadilla et al., 2013; Aggarwal et al., 2016) often rely on user behavioral data to learn user preferences to provide personalized services in multiple applications of World Wide Web. Given that user interests tend to evolve over time, to more accurately capture the short-term and immediate trend of user interest changes, sequential recommendation (SR) (Fang et al., 2020; Xie et al., 2022) has gradually become a key focus in research and applications. Different from traditional content-based (Pazzani and Billsus, 2007; Lops et al., 2011) or collaborative filtering RSs (He et al., 2017, 2020), SR emphasizes the temporal order and contextual information of behaviors, which uses the user’s past behavior sequence to model user interests and predict the items that may be of interest next. Most current SR methods model user interests as a single, comprehensive representation to capture the user’s overall interests. However, users may be interested in multiple types of content over a while, so there is a clear representational bottleneck in modeling users’ varying interests using a single-interest learning model.
In the MIND model (Li et al., 2019) proposed in 2019, the concept of multi-interest learning is introduced for the first time. The model employs the dynamic routing mechanism of capsule networks to capture the diverse interests of users. Since then, the research of multi-interest learning (Du et al., 2024; Liu et al., 2024; Yan et al., 2024; Pei et al., 2024) has received increasing attention. Current multi-interest approaches mainly rely on user behavioral data to learn users’ interests. However, most of this behavioral data consists of implicit feedback (Wang et al., 2021), such as clicking, browsing, and purchasing behaviors, which do not always accurately reflect the user’s true preferences. It may contain noisy data (Han et al., 2024; Qiao et al., 2023), such as records generated by mis-touching or non-true expressions of interest due to popularity bias, which may interfere with the model’s understanding of the user’s actual preferences. In addition, user behavioral data usually exhibits a high degree of sparsity, which further limits the model’s ability to capture users’ diverse and dynamically changing interests.
With the rapid development of natural language processing (NLP), the era of large language models (LLMs) (OpenAI, 2022; Gao et al., 2023) has arrived. LLMs’ outstanding language understanding, logical inference, and content generation capabilities provide new ideas and methods for solving the above challenges. That is, LLM is not only capable of extracting deep semantic information from text but also contains extensive knowledge of the outside world. With its superior contextual understanding, LLM effectively filters out irrelevant information and accurately captures user interests. Numerous studies (Cui et al., 2022; Geng et al., 2022; Li et al., 2023b; Zhang et al., 2023a; Yang et al., 2024) have shown that applying LLMs to RS demonstrates promising results in solving the cold-start problem and improving the interpretability of recommendation. However, since SR has high real-time requirements and the fine-tuning and inference of LLM requires a lot of time and computing resources, the fundamental problem of combining large models (LLMs) and small models (traditional algorithms) remains unsolved.
As mentioned above, multi-interest learning aims to learn the semantic-aware sub-interest, which happens to hold the same view as the LLM-enhanced RS. Therefore, combining the learned semantic information from LLMs with behavioral data is promising yet still very challenging: how to perform the alignment and how to ensure high efficiency?
To address the above challenges, we propose an LLM-assisted Explicit and Implicit Multi-interest Learning Framework (shorten as EIMF). Specifically, EIMF can be divided into two modules, the implicit behavioral interest module (IBIM) and the explicit semantic interest module (ESIM). In the training phase, IBIM follows the traditional multi-interest SR model, which can learn the implicit behavioral interests of users from their interaction data with items. ESIM, on the other hand, first analyzes user-generated text sequences using the clustering algorithm to identify representative samples. Next, these typical samples are subjected to deep inference using an LLM to reveal the users’ explicit semantic interests. Finally, we match semantic and behavioral interests based on “id-cluster numbering”, with joint learning on two auxiliary tasks (text classification and modal alignment), which can help fuse semantic information into behavioral interest representations. In the testing phase, accurate recommendations can be achieved based solely on user behavior data.
In summary, the contributions of this paper are as follows:
  • We answer the question of how to combine large&\&&small models in RS from the perspective of multi-interest learning and semantic alignment. We propose an LLM-assisted multi-interest learning framework, which can fully use the semantic reasoning capabilities of LLM and enrich the user’s behavioral learning, thereby significantly improving the accuracy of recommendation results.
  • We propose an efficient typical sample strategy to reduce the inference cost of LLM and enrich the user interest representation through multi-task learning in the training phase so that more accurate recommendations can be achieved in the testing phase by only using user behavior data.
@@
  • We conduct extensive experiments on real-world datasets, and the results show that our EIMF can significantly and stably improve the recommendation performance and exhibit excellent generalization capabilities.
@@ 2.1. Sequential Recommendation
Early SR research (Rendle et al., 2010; He and McAuley, 2016) focused on using the Markov Chain model to capture the transition probability between user behaviors to make recommendations. With the continuous advancement of deep learning technology, neural networks have become the mainstream method for SR. Among these methods, recurrent neural networks (RNNs) (Sherstinsky, 2020) are widely used due to their ability to effectively process sequence data. For example, GRU4Rec (Hidasi, 2015) uses GRU to capture changes in the user’s interest in the current session. Wu et al. (Wu et al., 2017) proposed RRN, which is implemented through a long short-term memory (LSTM) autoregressive model that captures dynamics and low-rank decomposition. With the emergence of the Transformer model (Vaswani, 2017), researchers have discovered that the attention mechanism can adaptively adjust the degree of attention to each element in the input information, achieving more accurate learning of user interests. NARM (Li et al., 2017) and SASRec (Kang and McAuley, 2018) use the attention mechanism to model the user’s sequential behavior, and capture the user’s main purpose. Bert4Rec (Sun et al., 2019) performs bidirectional encoding of user sequences based on the Bert structure and combines context to predict randomly masked items.
Although these SR models perform well in performance, they usually simplify user interests into a single embedding vector, resulting in recommendation results that cannot fully reflect the user’s actual multi-dimensional preferences, which in turn affects the personalization level and user experience.
@@ 2.2. Multi-interest Learning
As the research continues to deepen, researchers have gradually realized that users’ interests are often diverse, and a single interest modeling method makes it difficult to capture users’ complex and dynamically changing interests accurately. Alibaba’s research team proposed the MIND model (Li et al., 2019), which introduced the concept of multi-interest modeling. The model uses a dynamic routing mechanism to build a multi-interest extraction layer, which can effectively mine users’ multiple interests from their historical behavior data. The ComiRec (Cen et al., 2020) model explores multi-interest learning methods using dynamic routing and self-attention mechanisms. Tan et al. proposed the SINE model (Tan et al., 2021), in which the sparse interest module can adaptively infer the sparse concept set of each user from a large concept pool, and the interest aggregation module is used to model multiple user interests. Xie et al. (Xie et al., 2023) proposed the REMI framework, which uses an interest-aware hard negative mining strategy to effectively train discriminative representations and a routing regularization method to prevent interest routing collapse. Zhu et al. (Zhu et al., 2024) proposed the HPCL4SR model, in which category information was introduced into the model to construct a global graph to filter high-level preferences and used them as positive examples, and used contrastive learning to distinguish the differences between multiple interests based on user-item interaction information. PoMRec (Dong et al., 2024) first inserts specific prompts into user interactions to adapt them to the multi-interest extractor and aggregator and then utilizes the mean and variance embedding of user interactions to embed users’ multiple interests.
@@
2.3. LLM-assisted Recommender System
@@
As LLMs continue to demonstrate their superior capabilities, more and more research has begun to explore the application of LLMs in RS. LLM-assisted RS can analyze text information related to users and items, thereby understanding user interests more deeply. ChatRec (Gao et al., 2023) combines conversational AI, such as ChatGPT, with an RS that converts user profiles and historical interactions into prompts, relying only on contextual learning for effective recommendations without training. InstructRec (Zhang et al., 2023b) adapts to the recommendation task by way of LLM instruction tuning. By combining the strengths of traditional CTR models with pre-trained language models, the CTRL framework (Li et al., 2023a) aims to integrate and utilize information from different modalities more effectively. E4SRec solves the problem of representing ID information by injecting the ID embedding of items in the SR model into the LLM. SAID (Hu et al., 2024) uses a projector module to convert item IDs into embedding vectors, which are fed into LLM to obtain item embeddings containing fine-grained semantic information. RLMRec (Ren et al., 2024)incorporates auxiliary text signals, uses LLM for user/item analysis, and aligns the semantic space of LLM with collaborative relationship signals through cross-view alignment.
@@
Although LLM can significantly improve the performance of RS, LLM-assisted RS often faces challenges such as real-time requirements and large computational resources during training and deployment, which limits its feasibility for direct application in industrial environments.
Figure 1. EIMF overall architecture diagram.
3. Methodology
In Figure 1, we show the overall architecture of EIMF. In this section, we introduce the problem definition of SR in Section 3.1, then introduce EISM in Section 3.2, IBIM in Section 3.3, and finally introduce the training and serving of the framework in Section 3.4.

### v3

# LLM-based Bi-level Multi-interest Learning Framework for Sequential Recommendation

@@ Abstract.
Sequential recommendation (SR) has garnered significant attention due to its ability to capture users’ dynamic preferences. To further improve recommendation accuracy, recent research has introduced multi-interest learning into SR, enabling the modeling of user interests from multiple dimensions. However, most existing multi-interest SR approaches rely primarily on implicit feedback, which is inherently noisy and sparse, limiting the effectiveness of these models in delivering satisfactory recommendations. The rapid advancement of large language models (LLMs) has demonstrated their strong and robust reasoning capabilities, even on low-quality textual data, offering a promising avenue for uncovering users’ diverse interests. Nevertheless, the high computational cost and response latency of LLMs present substantial obstacles to their practical integration into SR systems.
To address these issues, we propose a novel LLM-based multi-interest learning framework for SR that captures user interests from both implicit behavioral and explicit semantic perspectives. The framework consists of two modules: the Implicit Behavioral Interest Module (IBIM) and the Explicit Semantic Interest Module (ESIM). IBIM employs a traditional multi-interest SR model to learn representations from user behavior data. ESIM first uses clustering to identify typical, information-rich samples, then applies prompt engineering to guide the LLM in semantically reasoning about these samples and extracting their underlying multi-interest representations. We then integrate the semantic insights from ESIM into the behavioral representations of IBIM through two carefully designed auxiliary tasks: modality alignment and semantic prediction. After that, only IBIM is responsible for inference, enabling efficient, LLM-free recommendations and avoiding latency issues. Extensive experiments on four real-world datasets demonstrate the effectiveness and practicality of our framework.
Recommender System; Sequential Recommendation; Large Language Model; Multi-interest Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
@@ 1. Introduction
In many real-world scenarios, such as e-commerce, social media, and online news platforms, user interests exhibit chronological evolution (Pazzani and Billsus, 2007; Bobadilla et al., 2013). As a result, Sequential Recommendation (SR) has emerged as an advanced and increasingly prominent subfield of Recommender Systems (RSs), attracting substantial attention in both academic research and practical applications (Fang et al., 2020; Xie et al., 2022). Unlike traditional RS approaches (Lops et al., 2011; Aggarwal et al., 2016; He et al., 2017, 2020), SR explicitly models the temporal dynamics and contextual dependencies of user behavior sequences, leveraging historical interactions to infer evolving user preferences and predict subsequent items of interest. Recent advances in SR have demonstrated that multi-interest learning significantly outperforms single-interest models (Hidasi, 2015; Kang and McAuley, 2018; Harte et al., 2023). By disentangling and capturing multiple latent preference factors, multi-interest methods offer a more expressive and fine-grained representation of user intent, thereby enhancing the system’s ability to model diverse behavioral patterns and make more accurate recommendations.
As research on multi-interest learning continues to gain momentum (Du et al., 2024; Liu et al., 2024; Yan et al., 2024; Pei et al., 2024), its inherent limitations are becoming increasingly evident. Most existing multi-interest approaches primarily depend on user implicit feedback data (e.g., clicks, views, and browsing histories) to infer underlying preferences. Mining multi-interests on such data is challenging due to the low-quality characteristics: (1) The implicit feedback data is ambiguous. For example, a user might click on a product out of curiosity or due to an eye-catching thumbnail, without any actual purchase intention. Similarly, prolonged viewing time may be caused by confusion or distraction rather than true interest. These ambiguities illustrate the inherent unreliability of implicit feedback in accurately capturing user intent. (2) The implicit feedback is frequently contaminated by noise (Han et al., 2024; Qiao et al., 2023), including accidental interactions or misleading signals introduced by popularity bias, where widely promoted or trending items attract interactions irrespective of a user’s true preference. Such noise may distort the model’s interpretation, leading to suboptimal or even irrelevant recommendations. (3) User behavior data is typically highly sparse, especially for users with limited activity histories (e.g., cold-start users or niche-interest users), posing further challenges to the accurate modeling of interest diversity and temporal dynamics. These limitations hinder the model’s ability to capture the evolving and multifaceted nature of user preferences comprehensively. Therefore, despite the notable advancements that multi-interest learning has brought to personalized recommendation, critical challenges remain, particularly in mitigating the effects of noisy feedback and overcoming the constraints imposed by data sparsity.
Recently, Large Language Models (LLMs) have achieved remarkable success in the field of Natural Language Processing (NLP), demonstrating exceptional capabilities in language understanding, logical reasoning, and content generation (OpenAI, 2022; Gao et al., 2023). Recognizing their potential, researchers have started to exploit LLMs’ powerful contextual comprehension to filter out irrelevant information, as well as their rich open-domain semantic knowledge to alleviate data sparsity, resulting in promising advances in recommendation tasks (Cui et al., 2022; Geng et al., 2022; Li et al., 2023b; Zhang et al., 2023a; Yang et al., 2024). Despite these encouraging developments, integrating LLMs into multi-interest SR remains a non-trivial task due to several critical challenges:
  • Limited understanding of multi-interest structures: Although LLMs exhibit strong language understanding, their capacity to identify and disentangle multiple user interests from sequential behavior remains underexplored.
  • Scalability concerns: The large-scale nature of user-item interaction data, combined with the high computational demands of LLMs, introduces substantial challenges in terms of efficiency and scalability when extracting user interests.
  • Latency constraints: Real-world RSs typically demand low-latency responses, making it difficult to fully leverage the power of LLMs without compromising serving efficiency.
Therefore, while LLMs hold great promise for enhancing multi-interest SR, realizing their full potential requires addressing these fundamental issues related to interest disentanglement, scalability, and serving latency.
In this paper, we propose an LLM-based Explicit and Implicit Multi-interest Learning Framework (EIMF) to investigate how large language models can be effectively and efficiently leveraged to enhance multi-interest learning in SR tasks. Specifically, EIMF comprises two core components: an Implicit Behavioral Interest Module (IBIM) and an Explicit Semantic Interest Module (ESIM). The IBIM is designed to model users’ implicit behavioral interests by capturing patterns from their historical interaction sequences using conventional multi-interest SR techniques. At the same time, the ESIM utilizes the contextual and semantic understanding capabilities of LLMs to extract users’ explicit semantic interests from textualized interaction histories. To address the limited understanding of multi-interest structures and scalability concerns, ESIM adopts a clustering algorithm to compress the user interaction texts and identify a set of typical samples. These representative samples are then used to guide the LLM in generating high-quality semantic interest representations, thereby reducing noise and enhancing interpretability. To address latency constraints, the explicit semantic information extracted by ESIM is incorporated into the behavioral representations generated by IBIM through the joint optimization of two auxiliary tasks: modality alignment and semantic prediction. These tasks enforce semantic consistency and mutual enhancement between the implicit and explicit interest spaces, thereby enabling efficient recommendations based solely on implicit behavioral representations during the service stage. By bridging behavioral signals and semantic understanding, EIMF provides a model-agnostic framework for capturing the multifaceted nature of user interests in SR. In summary, the contributions of this paper are as follows:
  • We propose an LLM-based multi-interest learning framework for SR, which uses the powerful reasoning ability of LLM to capture explicit multi-interest representations at the semantic level while modeling behavior representations with traditional SR models. By effectively combining semantic and behavioral information, the recommendation accuracy is significantly improved.
  • We reduce the amount of data required for LLM inference, thereby lowering resource costs, by selecting typical samples through clustering. Moreover, we enhance individual users’ interest representations by leveraging the enriched group features of these typical samples. In addition, the framework separates the training phase from the serving phase, with the LLM only playing a role during training, thereby ensuring the low-latency response of the RS in the serving phase.
@@
  • We conduct extensive experiments on four real-world datasets, and the results show that our EIMF can significantly and stably improve the recommendation performance and exhibit excellent generalization capabilities.
@@ 2.1. Sequential Recommendation
Early SR research (Rendle et al., 2010; He and McAuley, 2016) focused on using the Markov Chain model to capture the transition probability between user behaviors to make recommendations. With the continuous advancement of deep learning technology, neural networks have become the mainstream method for SR. Among these methods, Recurrent Neural Networks (RNNs) (Sherstinsky, 2020) are widely used due to their ability to effectively process sequence data. For example, GRU4Rec (Hidasi, 2015) uses a Gated Recurrent Unit (GRU) to capture changes in the user’s interest in the current session. Wu et al. (Wu et al., 2017) proposed RNN, which is implemented through a Long Short-Term Memory (LSTM) autoregressive model that captures dynamics and low-rank decomposition. With the emergence of the Transformer model (Vaswani, 2017), researchers have discovered that the attention mechanism can adaptively adjust the degree of attention to each element in the input information, achieving more accurate learning of user interests. NARM (Li et al., 2017) and SASRec (Kang and McAuley, 2018) use the attention mechanism to model the user’s sequential behavior, and capture the user’s main purpose. BERT4Rec (Sun et al., 2019) performs bidirectional encoding of user sequences based on the BERT structure and combines context to predict randomly masked items. S3-Rec (Zhou et al., 2020) is based on a self-attention neural structure and designs four auxiliary self-supervision objectives. It uses the mutual information maximization principle to learn the correlations between attributes, items, subsequences, and sequences to improve sequential recommendations. ICLRec (Chen et al., 2022) introduces a latent variable to represent user intent and learns its distribution function through clustering. It then leverages contrastive self-supervised learning to incorporate the learned intents into the SR model, maximizing the consistency between sequence views and their corresponding intents. PDRec (Ma et al., 2024) is a plug-and-play diffusion SR framework that models user preferences via time-interval diffusion and enhances recommendation performance through behavior reweighting, positive augmentation, and noise-free negative sampling.
Although these SR models have achieved good results in performance, they generally tend to represent the user’s interests as a comprehensive embedding. This single-interest learning method may ignore the subtle differences in user interests, resulting in recommendation results that cannot fully reflect the user’s multi-dimensional preferences, thereby affecting the personalization level and user experience.
@@ 2.2. Multi-interest Learning
As the research continues to deepen, researchers have gradually realized that users’ interests are often diverse, and a single interest modeling method makes it difficult to capture users’ complex and dynamically changing interests accurately. Alibaba’s research team proposed the MIND model (Li et al., 2019), which introduced the concept of multi-interest modeling. The model uses a dynamic routing mechanism to build a multi-interest extraction layer, which can effectively mine users’ multiple interests from their historical behavior data. The ComiRec (Cen et al., 2020) model explores multi-interest learning methods using dynamic routing and self-attention mechanisms. Tan et al. proposed the SINE model (Tan et al., 2021), in which the sparse interest module can adaptively infer the sparse concept set of each user from a large concept pool, and the interest aggregation module is used to model multiple user interests. Xie et al. (Xie et al., 2023) proposed the REMI framework, which uses an interest-aware hard negative mining strategy to effectively train discriminative representations and a routing regularization method to prevent interest routing collapse. Zhu et al. (Zhu et al., 2024) proposed the HPCL4SR model, in which category information was introduced into the model to construct a global graph to filter high-level preferences and used them as positive examples, and used contrastive learning to distinguish the differences between multiple interests based on user-item interaction information. PoMRec (Dong et al., 2024) first inserts specific prompts into user interactions to adapt them to the multi-interest extractor and aggregator, and then utilizes the mean and variance embedding of user interactions to embed users’ multiple interests. NP-Rec combines the Mamba sequence model with neural processes to learn the distribution of user preference functions. NP-Rec (Jiang et al., 2025) combines the Mamba sequence model with neural processes to learn the distribution of user preference functions. This design enables NP-Rec to not only capture multiple evolving interests but also provide uncertainty estimation, offering a more robust and adaptive recommendation strategy. DMI-GNN (Lv et al., 2025) integrates graph neural networks with dynamic multi-interest learning for session-based recommendation. It introduces a multi-position pattern learning strategy and a dynamic regularization method that adaptively adjusts the distinctiveness of interest representations according to the session length.
@@
2.3. LLM-based Recommender System
@@
As LLMs continue to demonstrate their superior capabilities, more and more research has begun to explore the application of LLMs in RS. LLM-based RS can analyze text information related to users and items, thereby understanding user interests more deeply. ChatRec (Gao et al., 2023) combines conversational AI, such as ChatGPT, with an RS that converts user profiles and historical interactions into prompts, relying only on contextual learning for effective recommendations without training. InstructRec (Zhang et al., 2023b) adapts to the recommendation task by way of LLM instruction tuning. By combining the strengths of traditional CTR models with pre-trained language models, the CTRL framework (Li et al., 2023a) aims to integrate and utilize information from different modalities more effectively. E4SRec solves the problem of representing ID information by injecting the ID embedding of items in the SR model into the LLM. SAID (Hu et al., 2024) uses a projector module to convert item IDs into embedding vectors, which are fed into LLM to obtain item embeddings containing fine-grained semantic information. RLMRec (Ren et al., 2024)incorporates auxiliary text signals, uses LLM for user/item analysis, and aligns the semantic space of LLM with collaborative relationship signals through cross-view alignment. PO4ISR (Sun et al., 2024) enhances next-item prediction in session-based recommendation by first generating an initial prompt to guide LLMs based on diverse user intents. It further introduces an iterative prompt optimization mechanism that enables self-reflection and continual refinement for improved prediction accuracy. LLMEmb (Liu et al., 2025) is a LLM-based item embedding generator that produces high-quality embeddings through supervised contrastive fine-tuning and recommendation adaptation training, significantly improving sequential recommendation performance and alleviating the long-tail problem.
@@
Although LLM can significantly improve the performance of RS, LLM-based RS often faces challenges such as real-time requirements and large computational resources during training and deployment, which limits its feasibility for direct application in industrial environments.
Table 1. Description of notations.
Notations Descriptions
$\mathcal{U}$ The set of all users
$\mathcal{I}$ The set of all items, defined as =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{N}\}$, where $N$ is the number of unique items in the dataset
$S_{u}$ The historical interaction sequence of a user $u$, represented as $S_{u}=[i_{1}^{u},i_{2}^{u},\dots,i_{L}^{u}]$, where $L$ is the maximum length of the sequence ordered by interaction time
$S_{u,k}$ The interaction sequence of user $u$ at the previous $k$ steps, represented as $S_{u,k}=[i_{1}^{u},i_{2}^{u},\dots,i_{k}^{u}]$
$i_{k+1}^{u}$ The item that user $u$ may click at the next step after timestamp $k$
$y_{k}$, ${{y_{k}}^{\text{t}}}$ The ground truth for the recommendation task and semantic prediction task for candidate item $i_{k}$
$\hat{y_{k}}$, $\hat{{y_{k}}}^{\text{t}}$ The predicted label for the recommendation task and semantic prediction task for candidate item $i_{k}$
$\mathbf{S}$, $\mathbf{R}$, $\mathbf{A}$ Similarity matrix, responsibility matrix, and availability matrix in AP clustering. $s(i,k)$, $r(i,k)$, $a(i,k)$ are elements of $\mathbf{S}$, $\mathbf{R}$, and $\mathbf{A}$
$T$ The text form of the user’s historical click sequence, represented as $T=[t_{1},t_{2},\dots,t_{L}]$
$C$ The category of clustering, represented as $C=\{c_{1},c_{2},\dots,c_{G}\}$, where $G$ is the number of clusters
$\mathbf{T}$, $\mathbf{t}$ Text embeddings obtained using a pre-trained language model, e.g., $\mathbf{T}$ for sequence embeddings, $\mathbf{t}_{\text{tar}}^{u}$, $\mathbf{t}_{k}^{u}$ for target and candidate item embeddings
$\mathbf{E}$, $\mathbf{e}$ ID embeddings obtained using a multi-interests SR model, e.g., $\mathbf{E}$ for sequence embeddings, $\mathbf{e}_{\text{tar}}$, $\mathbf{e}_{k}$ for target and candidate item embeddings
$M_{\text{im}}$, $M_{\text{ex}}$ Number of implicit and explicit interests learned
$\mathbf{H}_{\text{im}}$, $\mathbf{H}_{\text{ex}}$ Implicit behavioral interest representation ($\mathbf{H}_{\text{im}}\in\mathbb{R}^{M_{\text{im}}\times d}$) and explicit semantic interest representation ($\mathbf{H}_{\text{ex}}\in\mathbb{R}^{M_{\text{ex}}\times d_{t}}$)
$d$, $d_{t}$ Embedding dimension for ID embeddings and text embeddings, respectively
$\mathbf{W}$, $b$ Weight matrix and biases
$PLM$, $LLM$, $AP$, $MI$ Pre-trained language model, large language model, affinity propagation algorithm, and multi-interest SR model
$p$, $\alpha$, $\beta$, $\gamma$, $\tau$ Hyperparameters: preference in AP, weights of contrastive and cosine losses, weight of auxiliary tasks, and temperature in contrastive loss
$\mathcal{L}$, $\mathcal{L}_{A}$, $\mathcal{L}_{R}$, $\mathcal{L}_{S}$ Total loss function, modality alignment task loss, recommendation task loss, and semantic prediction task loss
Figure 1. Overall Framework of EIMF.
3. Our Approach: Explicit and Implicit Multi-interest Learning Framework
This section presents the technical details of our EIMF framework. First, we give a brief statement for the sequential recommendation in Section 3.1. Then, Section 3.2 provides an overview of EIMF, accompanied by the illustration in Figure 1. The two key components of EIMF are introduced in Sections 3.3 and 3.4, each described in detail with corresponding pseudocode in Algorithms 1 and 2. Section 3.5 discusses the training and deployment of EIMF.